Compute the additional Medicare tax for the following taxpayers. If required, round your answers to the nearest dollar.

a. Mario, who is single, earns wages of $516,600 in 2018.
His total Medicare tax is $___________?

b. George and Shirley are married and file a joint return in 2018. During the year, George earns wages of $198,900, and Shirley earns wages of $298,350.
Their total Medicare tax is $___________?

c. Simon has net investment income of $47,080 and MAGI of $235,400 and files as a single taxpayer.
Simon's additional Medicare tax is $___________?
Answer #1

a

Additional medicare = (516,600 - 200,000) X 0.9% = 2,849.40

b

Additional Medicare = (198,900 + 298,350 - 250,000) X 0.9% = 2,225.25

c

Additional medicare = (235,400 - 200,000) X 0.9% = 318.60